# Bandits with Preference Feedback: A Stackelberg Game Perspective

### v2

@@ Appendix C Proofs for Bandits with Preference Feedback
This section presents the proof of main results in Section 5, and our additional contributions in the kernelized Preference-based setting.
@@ C.1 Equivalence of Preference-based and Logistic Losses
=$\stackrel{{\scriptstyle b_{i}=0}}{{=}}\left\lVert f\right\rVert^{2}_{k}\leq B^{2}.$
@@
Proof of Corollary 5.
Consider the utility function $f$ and define $h({\bm{x}},{\bm{x}}^{\prime})\coloneqq f({\bm{x}})-f({\bm{x}}^{\prime})$. Then by Proposition 4, $h$ is in RKHS of $k^{\mathrm{D}}$ with a $k^{\mathrm{D}}$-norm bounded by $B$. We may estimate $h$ by minimizing ${\mathcal{L}}_{k^{\mathrm{D}}}^{\mathrm{L}}(\cdot;H_{t})$. Now invoking Theorem 2 with the dueling kernel we have,
@@ C.2 Proof of the Preference-based Regret Bound
Recall Corollary 5, which states
@@
Note that $\omega_{t}({\bm{x}},{\bm{x}}^{\prime})=\omega_{t}({\bm{x}}^{\prime},{\bm{x}})$ by the symmetry of the dueling kernel $k^{\mathrm{D}}$. Furthermore, recall the notation $h({\bm{x}},{\bm{x}})=f({\bm{x}})-f({\bm{x}})$.
Proof of Theorem 6.
Step 1: First, we connect the term of ${\bm{x}}_{t}^{\prime}$ in the dueling regret defined in Equation 1 to that of ${\bm{x}}_{t}$. Note that both $s(f({\bm{x}}^{*})-f({\bm{x}}_{t}^{\prime}))$ and $s(f({\bm{x}}^{*})-f({\bm{x}}_{t}))$ are greater than $0.5$ due to the optimality of ${\bm{x}}^{\star}$ and the sigmoid function $s$ is concave on the interval $[0.5,\infty)$. Using the definition of concavity, we get
Using Equation 15, we can upper bound the dueling regret in Equation 1 as
First, consider the term $s(f({\bm{x}}^{\star})-f({\bm{x}}_{t}))$
$s(f({\bm{x}}^{\star})-f({\bm{x}}_{t}))-0.5$ $\leq s(h_{t}({\bm{x}}^{\star},{\bm{x}}_{t}))+\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\star})-0.5$
$\leq 0.5-s(h_{t}({\bm{x}}_{t},{\bm{x}}^{\star}))+\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\star})$ Sigmoid equality
$\leq 2\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\star})$ (17)
In the last inequality, we used that ${\bm{x}}_{t}\in{\mathcal{M}}_{t}$ implying that $0.5-s(h_{t}({\bm{x}}_{t},{\bm{x}}^{\star}))\leq\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\star})$. Combining Equation 16 and Equation 17 implies that
$2r_{t}^{D}\leq 4\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\star})+\frac{h({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})}{2}s(f({\bm{x}}^{\star})-f({\bm{x}}_{t}))$ (18)
Now, we bound $\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\star})$ by $\omega_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})$. If ${\bm{x}}_{t}={\bm{x}}^{\star}$, then $\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\star})=0\leq\omega_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})$. If ${\bm{x}}_{t}^{\prime}={\bm{x}}^{\star}$, then the two expressions are equivalent. Now, assume that ${\bm{x}}_{t}\not={\bm{x}}^{*}$ and ${\bm{x}}_{t}^{\prime}\not={\bm{x}}^{\star}$ and consider $\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\star})$.
Case 1: Assume that $\mathrm{UCB}_{t}({\bm{x}}_{t},{\bm{x}}^{\star})\leq\mathrm{UCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})$. Then,
$2\omega_{t}({\bm{x}}_{t},{\bm{x}}^{*})$ $=\mathrm{UCB}_{t}({\bm{x}}_{t},{\bm{x}}^{\star})-\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}^{\star})$
$\leq\mathrm{UCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})-\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}^{\star})$
$\leq\mathrm{UCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})-\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})=2\omega_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})$ (19)
where we used the assumption in the first inequality and the definition of ${\bm{x}}_{t}^{\prime}$ in the second inequality.
Case 2: Assume that $\mathrm{UCB}_{t}({\bm{x}}_{t},{\bm{x}}^{\star})\geq\mathrm{UCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})$. First note the following connection between $\mathrm{LCB}_{t}$ and $\mathrm{UCB}_{t}$.
$\mathrm{LCB}_{t}({\bm{x}},{\bm{x}}^{\prime})$ $=s(h_{t}({\bm{x}},{\bm{x}}^{\prime}))-\omega_{t}({\bm{x}},{\bm{x}}^{\prime})$
$=1-s(h_{t}({\bm{x}}^{\prime},{\bm{x}}))-\omega_{t}({\bm{x}}^{\prime},{\bm{x}})$
$=1-\mathrm{UCB}_{t}({\bm{x}}^{\prime},{\bm{x}})$ (20)
where we used the equality $s(z)=1-s(-z)$ in the second equality. Using Equation 20, the assumption of Case 2 can be rewritten as $\mathrm{UCB}_{t}({\bm{x}}_{t},{\bm{x}}^{\star})\geq\mathrm{UCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})$, implying that
$\mathrm{LCB}_{t}({\bm{x}}_{t}^{\prime},{\bm{x}}_{t})\geq\mathrm{LCB}_{t}({\bm{x}}^{\star},{\bm{x}}_{t}).$ (21)
Similarly, $\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})\leq\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}^{\star})$ implies
$\mathrm{UCB}_{t}({\bm{x}}_{t}^{\prime},{\bm{x}}_{t})\geq\mathrm{UCB}_{t}({\bm{x}}^{\star},{\bm{x}}_{t})$ (22)
Combining Equation 21 and Equation 22, we get
$2\omega_{t}({\bm{x}}_{t},{\bm{x}}^{*})$ $=\mathrm{UCB}_{t}({\bm{x}}^{\star},{\bm{x}}_{t})-\mathrm{LCB}_{t}({\bm{x}}^{\star},{\bm{x}}_{t})$
$\leq\mathrm{UCB}_{t}({\bm{x}}^{\star},{\bm{x}}_{t})-\mathrm{LCB}_{t}({\bm{x}}_{t}^{\prime},{\bm{x}}_{t})$
$\leq\mathrm{UCB}_{t}({\bm{x}}_{t}^{\prime},{\bm{x}}_{t})-\mathrm{LCB}_{t}({\bm{x}}_{t}^{\prime},{\bm{x}}_{t})=2\omega_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})$ (23)
Combining Equation 19 and Equation 23, we can rewrite the first term in Equation 18 to get
It remains to bound the second term of Equation 24. By the Mean-Value Theorem, $\exists z\in[0,h({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})]$ such that
Now since $\kappa=\sup_{z\leq B}1/\dot{s}(z)$ then,
Next, we consider the term $s(h({\bm{x}}_{t},{\bm{x}}_{t}^{\prime}))-0.5$ in Equation 25. Note that ${\bm{x}}_{t},{\bm{x}}_{t}^{\prime}\in{\mathcal{M}}_{t}$ implies that
@@
From Equation 26 and Equation 27, it follows that
$|s(h_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime}))-0.5|\leq\omega_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})$
@@
From Equation 28 and Equation 29, it follows that
@@
Combining Equation 30 with Equation 25, it follows
@@
Using Equation 31 in Equation 24 and the fact that $s(f({\bm{x}}^{*})-f({\bm{x}}_{t}))\leq 1$, we get
@@ C.3 Extending Algorithms for Linear Dueling Bandits to Kernelized Setting
Maximum Informative Pair Algorithm. Proposed in Saha [2021] for linear utilities, the MaxInP algorithm similarly maintains a set of plausible maximizer arms, and picks the pair of actions that have the largest joint uncertainty, and therefore are expected to be informative. Algorithm 3 present the kernelized variant of this algorithm. Using Corollary 5, we can show that the kernelized MaxInP also satisfies a $\tilde{\mathcal{O}}(\gamma_{T}\sqrt{T})$ regret.
@@
Let $\delta\in(0,1]$ and choose the exploration coefficient $\beta^{\mathrm{D}}_{t}(\delta)$ as defined in Corollary 5. Then MaxInP satisfies the anytime dueling regret guarantee of
@@
Proof of Theorem 15.
When selecting $({\bm{x}}_{t},{\bm{x}}^{\prime}_{t})$ according to Algorithm 3, we choose the pair via
@@
$\begin{split}s(h_{t}({\bm{x}}^{\star},{\bm{x}}_{t}))&\leq 1/2+\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\star})\\ s(h_{t}({\bm{x}}^{\star},{\bm{x}}^{\prime}_{t}))&\leq 1/2+\omega_{t}({\bm{x}}^{\prime}_{t},{\bm{x}}^{\star})\\ \end{split}$ (33)
where we have used the identity $s(-z)=1-s(z)$. Simultaneously for all $t\geq 1$, we can bound the single-step dueling regret with probability greater than $1-\delta$
@@
where for the first inequality we have invoked Corollary 5. Then for the regret satisfies
@@
and show its uniform validity in Lemma 17. We can now propose the Kernelized Logistic IDS algorithm with preference feedback in Algorithm 4, as a variant of the algorithm of Kirschner and Krause.
@@
Let $\delta\in(0,1]$ and for all $t\geq 1$, set the exploration coefficient as $\tilde{\beta}_{t}=\beta_{t}^{\mathrm{D}}(\delta)/L$. Then Algorithm 4 satisfies the anytime cumulative dueling regret guarantee of
@@
Initialize Set $(\beta_{t})_{t\geq 1}$ according to Theorem 2.
@@
Proof of Theorem 16.
Our approach closely follows the proof of Kirschner and Krause [2021, Theorem 1]. Let ${\mathcal{P}}(\cdot)$ show the set of continuous probability distributions over a domain. Define the expected average gap for a policy $\mu\in{\mathcal{P}}({\mathcal{X}}\times{\mathcal{X}})$
@@
Algorithm 4 draws actions via $\mu_{t}=(1-p_{t})\delta_{({\bm{x}}^{(1)}_{t},{\bm{x}}^{(1)}_{t})}+p_{t}\delta_{({\bm{x}}^{(1)}_{t},{\bm{x}}^{(2)}_{t})}$, where $\delta_{({\bm{x}},{\bm{x}}^{\prime})}$ denotes a Direct delta. Then by Kirschner et al. [2020, Lemma 1],
@@
since $s(\cdot)$ with its domain restricted to $[-2B,2B]$ is $L$-Lipschitz. It remains to bound $\Xi_{t}(\mu_{t})$, the expected information ratio for Algorithm 4. Now by definition of $\mu_{t}$
@@
where last inequality holds due to the following argument. Recall that $k({\bm{x}},{\bm{x}})\leq 1$, implying that $\sigma_{t}^{\mathrm{D}}({\bm{x}},{\bm{x}}^{\prime})^{2}\leq 4$ and therefore $\log(1+\sigma_{t}^{\mathrm{D}}({\bm{x}},{\bm{x}}^{\prime})^{2})\geq\log(1+(\lambda\kappa)^{-1})\sigma_{t}^{\mathrm{D}}({\bm{x}},{\bm{x}}^{\prime})^{2}/4$, similar to Lemma 14. To conclude the proof, from (35) and (36) it holds that
@@
C.4 Helper Lemmas for Section C.3
@@
Let $10<\delta<$ 1 and $f\in{\mathcal{H}}_{k}$. Suppose $\sup_{a\leq B}\dot{s}(a)=L$ and $\sup_{a\leq B}1/\dot{s}(a)=\kappa$. Then
@@
Assume $f\in{\mathcal{H}}_{k}$. Suppose $\sup_{a\leq B}1/\dot{s}(a)=\kappa$. Then for any $10<\delta<$ 1
@@
Proof of Lemma 18.
This lemma is effectively a weaker parallel of Corollary 5. We have
@@ Appendix D Details of Experiments
Figure 3: Confidence sets for an illustrative problem with $3$ arms at a single time step. Annotated arrows highlight the action selection for three common approaches. MaxMinLCB selects the action pair $(1,2)$ with the least regret. Upper-bound maximization (Optimism) and information maximization (Max Info) choose sub-optimal arms.
Test Environments. We use a wide range of target functions common to the optimization literature [Jamil and Yang, 2013], to evaluate the robustness of MaxMinLCB. The results are reported in Table 1 and Table 2. Note that for the experiments we negate them all to get utilities. We use a uniform grid of $100$ points over their specified domains and scale the utility values to $[-3,3]$.
@@
Algorithm 5 Doubler [Ailon et al., 2014]
@@
Algorithm 6 MultiSBM [Ailon et al., 2014]
@@
Algorithm 7 RUCB [Zoghi et al., 2014a]
@@
Hyper-parameters for Logistic Bandits. We set $\delta=0.1$ for all algorithms. For GP-UCB and LGP-UCB, we set $\beta=1$, and $0.25$ for the noise variance. We use the Radial Basis Function (RBF) kernel and choose the variance and length scale parameters from $[0.1,1.0]$ to optimize their performance separately. For LGP-UCB, we tuned $\lambda$, the $L2$ penalty coefficient in Proposition 1, on the grid $[0.0,0.1,1.0,5.0]$ and $B$ on $[1.0,2.0,3.0]$. The hyper-parameter selections were done for each algorithm separately to create a fair comparison.
@@
Pseudo-code for Baselines. Algorithm 5, Algorithm 6, and Algorithm 7 described the baselines used for the benchmark of Section 6.2. MaxInP and IDS are defined in Algorithm 3 and Algorithm 4, respectively, in Section C.3 alongside with their theoretical analysis. We note that Doubler includes an internal for-loop, therefore, we adjusted the time-horizon $T$ such that it observes the same number of feedback $y_{t}$ as the other algorithms for a fair comparison.
Computational Resources and Costs. We ran our experiments on a shared cluster equipped with various NVIDIA GPUs and AMD EPYC CPUs. Our default configuration for all experiments was a single GPU with 24 GB of memory, 16 CPU cores, and 16 GB of RAM. Each experiment of the $11$ configurations reported in Section 6.2 ran for about $12$ hours and the experiment reported in Section 6.1 ran for 5 hours. The total computational cost to reproduce our results is around 140 hours of the default configuration. Our total computational costs including the failed experiments are estimated to be 2-3 times more.

### v3

@@ Appendix C Proofs for Bandits with Preference Feedback
This section presents the proof of main results in Section˜5, and our additional contributions in the kernelized Preference-based setting.
@@ C.1 Equivalence of Preference-based and Logistic Losses
$\stackrel{{\scriptstyle b_{i}=0}}{{=}}\left\lVert f\right\rVert^{2}_{k}\leq B^{2}.$
@@
Proof of Corollary˜5.
Consider the utility function $f$ and define $h({\bm{x}},{\bm{x}}^{\prime})\coloneqq f({\bm{x}})-f({\bm{x}}^{\prime})$. Then by Proposition˜4, $h$ is in RKHS of $k^{\mathrm{D}}$ with a $k^{\mathrm{D}}$-norm bounded by $B$. We may estimate $h$ by minimizing ${\mathcal{L}}_{k^{\mathrm{D}}}^{\mathrm{L}}(\cdot;H_{t})$. Now invoking Theorem˜2 with the dueling kernel we have,
@@ C.2 Proof of the Preference-based Regret Bound
Recall Corollary˜5, which states
@@
Note that $\omega_{t}({\bm{x}},{\bm{x}}^{\prime})=\omega_{t}({\bm{x}}^{\prime},{\bm{x}})$ by the symmetry of the dueling kernel $k^{\mathrm{D}}$. Furthermore, recall the notation $h({\bm{x}},{\bm{x}}^{\prime})=f({\bm{x}})-f({\bm{x}}^{\prime})$.
Before providing the proof of Theorem˜6, we derive two important inequalities. Note that ${\bm{x}}_{t},{\bm{x}}_{t}^{\prime}\in{\mathcal{M}}_{t}$ implies that
@@
From Equation˜15 and Equation˜16, it follows that
$|s(h_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime}))-0.5|\leq\omega_{t}({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})$ (17)
@@
From Equation˜18 and Equation˜19, it follows that
@@
We are going to refer to Equation˜17 and Equation˜20 in the following proof.
Proof of Theorem˜6.
Step 1: First, we connect the term of ${\bm{x}}_{t}^{\prime}$ in the dueling regret defined in Equation˜1 to that of ${\bm{x}}_{t}$. Note that both $s(f({\bm{x}}^{*})-f({\bm{x}}_{t}^{\prime}))$ and $s(f({\bm{x}}^{*})-f({\bm{x}}_{t}))$ are greater than $0.5$ due to the optimality of $\bm{x}$$^{\star}$ and the sigmoid function $s$ is concave on the interval $0.5,\infty)$. Using the definition of concavity, we get
Using Equation˜21, we can upper bound the dueling regret in Equation˜1 as
First, consider the term $s(f({\bm{x}}^{\star})-f({\bm{x}}_{t}))$
$s(f({\bm{x}}^{\star})-f({\bm{x}}_{t}))-0.5$ $=0.5-s(f({\bm{x}}_{t})-f({\bm{x}}^{\star}))$ Sigmoid definition
$\leq 0.5-\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}^{\star})$
$\leq 0.5-\mathrm{LCB}_{t}({\bm{x}}_{t},{\bm{x}}^{\prime}_{t})$ Def. of ${\bm{x}}^{\prime}_{t}$
$\leq|0.5-s(h_{t}({\bm{x}}_{t},{\bm{x}}^{\prime}_{t}))|+\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\prime}_{t})$ Def. of $\mathrm{LCB}_{t}$
$\leq 2\omega_{t}({\bm{x}}_{t},{\bm{x}}^{\prime}_{t})$ (23)
Using Equation˜23, we can rewrite the first term in Equation˜22 to get
It remains to bound the second term of Equation˜24. By the Mean-Value Theorem, $\exists z\in[0,h({\bm{x}}_{t},{\bm{x}}_{t}^{\prime})]$ such that
Now since $\kappa=\sup_{z\leq B}1/\dot{s}(z)$ then,
Combining Equation˜20 with Equation˜25, it follows
@@
Using Equation˜26 in Equation˜24 and the fact that $s(f({\bm{x}}^{*})-f({\bm{x}}_{t}))\leq 1$, we get
@@ C.3 Extending Algorithms for Linear Dueling Bandits to Kernelized Setting
Maximum Informative Pair Algorithm. Proposed in saha2021optimal for linear utilities, the MaxInP algorithm similarly maintains a set of plausible maximizer arms, and picks the pair of actions that have the largest joint uncertainty, and therefore are expected to be informative. Algorithm˜3 present the kernelized variant of this algorithm. Using Corollary˜5, we can show that the kernelized MaxInP also satisfies a $\tilde{\mathcal{O}}(\gamma_{T}\sqrt{T})$ regret.
@@
Let $\delta\in(0,1]$ and choose the exploration coefficient $\beta^{\mathrm{D}}_{t}(\delta)$ as defined in Corollary˜5. Then MaxInP satisfies the anytime dueling regret guarantee of
@@
Proof of Theorem˜15.
When selecting $({\bm{x}}_{t},{\bm{x}}^{\prime}_{t})$ according to Algorithm˜3, we choose the pair via
@@
[…]
where we have used the identity s(z)=1s(z)s(-z)=1-s(z). Simultaneously for all $t\geq 1$, we can bound the single-step dueling regret with probability greater than $1-\delta$
@@
where for the first inequality we have invoked Corollary˜5. Then for the regret satisfies
@@
and show its uniform validity in Lemma˜17. We can now propose the Kernelized Logistic IDS algorithm with preference feedback in Algorithm˜4, as a variant of the algorithm of kirschner2021bias.
@@
Let $\delta\in(0,1]$ and for all $t\geq 1$, set the exploration coefficient as $\tilde{\beta}_{t}=\beta_{t}^{\mathrm{D}}(\delta)/L$. Then Algorithm˜4 satisfies the anytime cumulative dueling regret guarantee of
@@
Initialize Set $(\beta_{t})_{t\geq 1}$ according to Theorem˜2.
@@
Proof of Theorem˜16.
Our approach closely follows the proof of kirschner2021bias. Let ${\mathcal{P}}(\cdot)$ show the set of continuous probability distributions over a domain. Define the expected average gap for a policy $\mu\in{\mathcal{P}}({\mathcal{X}}\times{\mathcal{X}})$
@@
Algorithm˜4 draws actions via $\mu_{t}=(1-p_{t})\delta_{({\bm{x}}^{(1)}_{t},{\bm{x}}^{(1)}_{t})}+p_{t}\delta_{({\bm{x}}^{(1)}_{t},{\bm{x}}^{(2)}_{t})}$, where $\delta_{({\bm{x}},{\bm{x}}^{\prime})}$ denotes the Dirac delta. Then by kirschner2020information,
@@
since $s(\cdot)$ with its domain restricted to [2B,2B][-2B,2B] is $L$-Lipschitz. It remains to bound $\Xi_{t}(\mu_{t})$, the expected information ratio for Algorithm˜4. Now by definition of $\mu_{t}$
@@
where last inequality holds due to the following argument. Recall that $k({\bm{x}},{\bm{x}})\leq 1$, implying that $\sigma_{t}^{\mathrm{D}}({\bm{x}},{\bm{x}}^{\prime})^{2}\leq 4$ and therefore $\log(1+\sigma_{t}^{\mathrm{D}}({\bm{x}},{\bm{x}}^{\prime})^{2})\geq\log(1+(\lambda\kappa)^{-1})\sigma_{t}^{\mathrm{D}}({\bm{x}},{\bm{x}}^{\prime})^{2}/4$, similar to Lemma˜14. To conclude the proof, from (30) and (31) it holds that
@@
C.4 Helper Lemmas for Section˜C.3
@@
Let $0<\delta<1$ and $f\in{\mathcal{H}}_{k}$. Suppose $\sup_{a\leq B}\dot{s}(a)=L$ and $\sup_{a\leq B}1/\dot{s}(a)=\kappa$. Then
@@
Assume $f\in{\mathcal{H}}_{k}$. Suppose $\sup_{a\leq B}1/\dot{s}(a)=\kappa$. Then for any $0<\delta<1$
@@
Proof of Lemma˜18.
This lemma is effectively a weaker parallel of Corollary˜5. We have
@@ Appendix D Details of Experiments
Figure 3: Confidence sets for an illustrative problem with 33 arms at a single time step. Annotated arrows highlight the action selection for three common approaches. MaxMinLCB selects the action pair $(1,2)$ with the least regret. Upper-bound maximization (Optimism) and information maximization (Max Info) choose sub-optimal arms.
Test Environments. We use a wide range of target functions common to the optimization literature [jamil2013literature], to evaluate the robustness of MaxMinLCB. The results are reported in Table˜1 and Table˜2. Note that for the experiments we negate them all to get utilities. We use a uniform grid of $100$ points over their specified domains and scale the utility values to [3,3][-3,3].
@@
Algorithm 5 Doubler [ailon2014reducing]
@@
Algorithm 6 MultiSBM [ailon2014reducing]
@@
Algorithm 7 RUCB [zoghi2014relative]
@@
Hyper-parameters for Logistic Bandits. We set $\delta=0.1$ for all algorithms. For GP-UCB and LGP-UCB, we set $\beta=1$, and $0.25$ for the noise variance. We use the Radial Basis Function (RBF) kernel and choose the variance and length scale parameters from $[0.1,1.0]$ to optimize their performance separately. For LGP-UCB, we tuned $\lambda$, the $L2$ penalty coefficient in Proposition˜1, on the grid $[0.0,0.1,1.0,5.0]$ and $B$ on $[1.0,2.0,3.0]$. The hyper-parameter selections were done for each algorithm separately to create a fair comparison.
@@
Pseudo-code for Baselines. Algorithm˜5, Algorithm˜6, and Algorithm˜7 described the baselines used for the benchmark of Section˜6.2. MaxInP and IDS are defined in Algorithm˜3 and Algorithm˜4, respectively, in Section˜C.3 alongside with their theoretical analysis. We note that Doubler includes an internal for-loop, therefore, we adjusted the time-horizon $T$ such that it observes the same number of feedback $y_{t}$ as the other algorithms for a fair comparison.
Computational Resources and Costs. We ran our experiments on a shared cluster equipped with various NVIDIA GPUs and AMD EPYC CPUs. Our default configuration for all experiments was a single GPU with 24 GB of memory, 16 CPU cores, and 16 GB of RAM. Each experiment of the $11$ configurations reported in Section˜6.2 ran for about $12$ hours and the experiment reported in Section˜6.1 ran for 5 hours. The total computational cost to reproduce our results is around 140 hours of the default configuration. Our total computational costs including the failed experiments are estimated to be 2-3 times more.